OpenStudy (anonymous):

solve the equation , 7+6n=-5

OpenStudy (anonymous):

7+6n=-5 -7 -7 6n=-12 n=-2

OpenStudy (anonymous):

can you help me also solve 6b+4=2(2b+1)

OpenStudy (anonymous):

6b+4=2(2b+1) 6b+4=4b+2 -4b -4b 2b+4=2 -4 -4 2b=-2 b=-1